OpenStudy (anonymous):

im need help trying to solve: ?(3.8)=9(4)-9(?)

OpenStudy (anonymous):

?(3.8) + 9(?)=9(4)-9(?) + 9(?) .... I added 9(?) to both sides ?(3.8+9) = 36 .... multiplied 9 and 4 and added everything related to ? ?(13.7) / 13.7 = 36 / 13.7 ?=2.627737